# Dynamic Matroids: Base Packing and Covering

### v1

@@ 1 Introduction
One well-studied example of a matroid is the graphic matroid, where $E$ is some set of edges and $A\subseteq E$ is independent if and only if it is acyclic. Base packing corresponds to packing disjoint spanning trees and the base covering number corresponds to the arboricity of the graph.
@@
For dynamic matroids, we consider element insertions and deletions to the ground set. [^2] The decremental version of this has been studied in [BlikstadMNT23], and we argue below that this fully dynamic version is the natural and most general definition. When an element $e$ is deleted, the independent sets $\mathcal{I}$ simply restrict to the sets without $e$: $\{I\in\mathcal{I}:e\notin I\}$. In other words, $\mathcal{M}$ is restricted to $E\setminus\{e\}$. When an element $E$ is inserted, the adversary also decides on a collection $\mathcal{I}_{e}$ such that $(E\cup\{e\},\mathcal{I}\cup\mathcal{I}_{e})$ is a matroid and $e\in I$ for every $I\in\mathcal{I}_{e}$. This means that inserting and then deleting the same element results in the same matroid. The algorithm receives the element updates and can query the (new) independence sets.
@@
Dynamic Graph Algorithms.
De Vos and Christiansen [deVosC24] give fully dynamic algorithms for $(1\pm\varepsilon)$-approximate arboricity with $O(\operatorname{poly}(\log|V|,\varepsilon^{-1}))$ update time against an adaptive adversary. Banerjee, Raman, and Saurabh [BanerjeeRS20] maintain the exact arboricity with $\tilde{O}(|E|)$ update time.
Dynamic tree-packing has been implicitly studied by Thorup [Thorup07]. This paper uses dynamic tree-packing due to its relation to min-cut. However, it implies a $(1\pm\varepsilon)$-approximation of the fractional tree packing number [^4] We realize that the naming conventions here overlap. Rather than disregarding the conventions completely, we write ‘tree-packing’ for the collection of trees and ‘tree packing’ when talking about the tree packing number. $\Phi\leq\Phi_{\max}$ in $\tilde{O}(\Phi_{\max}^{2}\log^{6}|V|/\varepsilon^{-4})$ amortized update time.
@@
There exists a deterministic algorithm that, given a dynamic matroid $\mathcal{M}$ with weight function $w\colon E\to[1,\dots,W]$, maintains a minimum weight base, where each update uses $O(\log(Wn))$ rank-queries.
Blikstad, Mukhopadhyay, Nanongkai, and Tu [BlikstadMNT23] provide an algorithm for maintaining a minimum weight basis under deletions. Each update requires $\tilde{O}(\sqrt{\operatorname{rk}(E)})$ worst-case rank-queries. However, using a ‘dynamic rank oracle’, they also ensure $\tilde{O}(\sqrt{\operatorname{rk}(E)})$ worst-case update time. The algorithm is based on the MST algorithm with $\tilde{O}(\sqrt{|V|})$ update time by Frederickson [Frederickson85], combined with the sparsification technique of Eppstein, Galil, Italiano, and Nissenzweig [EppsteinGIN97]. It seems likely that such a result can be extended to the fully dynamic setting. Blikstad et al. [BlikstadMNT23] use this as a subroutine and the authors seem to have optimized their result for their application in solving matroid union. The difference to our result, Proposition˜1.1, is that we have a much lower query time, but do not give guarantees on the update time.
Both algorithms use a rank oracle. Sometimes, algorithms are developed using an independence oracle, which only provides whether $A\in\mathcal{I}$. Clearly, the rank oracle is stronger than the independence oracle. Statically, it is even known that it is strictly stronger: computing a minimum weight base can be done with $n$ simultaneous rank-queries, while this needs $\tilde{\Omega}(n^{1/3})$ rounds of simultaneous independence-queries [KarpUW88]. It would be interesting to see if there exists a decremental algorithm for maintaining a minimum weight base using a sublinear number of independence queries.
@@
The integral packing number is $\lfloor\Phi_{\mathcal{M}}\rfloor$. Edmonds [Edmonds65] showed that $\lfloor\Phi_{\mathcal{M}}\rfloor$ equals the number of disjoint bases that can be packed in $\mathcal{M}$.
@@
We define the integral covering number as $\lceil\beta_{\mathcal{M}}\rceil$. Edmonds [Edmonds65partition] showed that $\lceil\beta_{\mathcal{M}}\rceil$ equals the minimum number of bases necessary to cover $\mathcal{M}$. We omit the subscript if the matroid is clear from the context.
In this paper, we give the first dynamic algorithms for approximating the fractional packing number and the fractional covering number.
@@
Let $\varepsilon\in(0,1/2)$ be a parameter and let $\mathcal{M}$ be a dynamic matroid that at any moment contains at most $n$ elements. If $\Phi$ is bounded by $\Phi_{\max}$, we can deterministically maintain a $(1\pm\varepsilon)$-approximation of the fractional packing number with $O(\Phi_{\max}^{2}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ worst-case rank-queries per update or $O(\Phi_{\max}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ amortized rank-queries per update.
Since we give the first dynamic algorithm, our result can only be compared to using the deterministic static algorithm from Chekuri and Quanrud [ChekuriQ17] to recompute the fractional packing number from scratch after every update. This would take $\tilde{O}(n\Phi_{\max}/\varepsilon^{2})$ independence-queries per update. Hence, our algorithm provides an exponential improvement in terms of $n$.
@@
Let $\varepsilon\in(0,1/2)$ be a parameter and let $\mathcal{M}$ be a dynamic matroid that at any moment contains at most $n$ elements. If $\beta$ is bounded by $\beta_{\max}$, we can deterministically maintain a $(1\pm\varepsilon)$-approximation of the fractional covering number with $O(\beta_{\max}^{2}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ worst-case rank-queries per update.
Again, there are no preexisting dynamic algorithms to compare our result to. Even static deterministic base covering is not as well studied. Quanrud [Quanrud24] conjectures that the techniques from Chekuri and Quanrud [ChekuriQ17] extend to deterministic approximate base covering in $\tilde{O}(n\beta_{\max}/\varepsilon^{2})$ independence-queries, which would transfer to $\tilde{O}(n\beta_{\max}/\varepsilon^{2})$ independence-queries per update.
@@
Once again, we compare to recomputing from scratch after every update using a static algorithm and obtain an exponential improvement. The state of the art for computing the static fractional covering number is by Quanrud [Quanrud24] and would result in $\tilde{O}(n+\operatorname{rk}(E)/\varepsilon^{3})$ independence-queries per update.
@@ 1.1 Technical Overview
Our main technical tool are base collections, which we will introduce first. Then, we will sketch how use them to obtain our packing and covering results, Theorems˜1.2, 1.3 and 1.4.
1.1.1 Base Collections.
@@
Next, we want to define some ‘ideal’ relative load. Hereto, we first prove a corollary using Edmond’s theorem on base packing (Theorem˜A.1). We show that the lowest possible maximum load in base collections relate to the base packing number.
@@
We show that this is well-defined in Appendix˜A.
@@
In the graphic matroid, it has been shown [Thorup07] that a greedy tree-packing $\mathcal{T}$ of size $\Theta(\Phi\log|V|/\varepsilon^{2})$ approximates the ideal packing well since
@@
for all $e\in E$. We show a stronger statement for general matroids, for which we introduce the parameter $\gamma$ to provide a better trade-offs between the error and the number of greedy bases.
@@
The improvement over [Thorup07] in the special case of the graphic matroid can be seen as follows: Equation˜1 can be reformulated to obtain $|\ell^{\mathcal{B}}(e)-\ell^{*}(e)|\leq\varepsilon/\gamma$. However, it needs $|\mathcal{B}|=\Theta\left(\tfrac{\gamma^{2}}{\Phi}\log n/\varepsilon^{2}\right)$ bases. That is, it has a quadratic dependence on $\gamma$, where we obtain a linear dependence.
In other words, Thorup [Thorup07] showed that the values are an additive approximation, since the error is independent of $\ell^{*}(e)$. We show a multiplicative approximation, since the error depends linearly on $\ell^{*}(e)$. For our application, the latter leads to stronger results. To be precise, Lemma˜1.6 shows that a collection of at least $3\beta\log n/\varepsilon^{2}$ bases directly gives a $(1\pm\varepsilon)$-approximation of the covering number. Using Thorup [Thorup07] gives $|\mathcal{B}|=\Theta\left(\tfrac{\beta^{2}}{\Phi}\log n/\varepsilon^{2}\right)$.
@@
To the best of our knowledge, the community was not aware of this property for the graphic matroid either. In particular, it means that we simplify the result of de Vos and Christiansen [deVosC24] for arboricity. They introduce an intricate procedure to artificially maintain $\Phi\approx\beta$ by adding virtual edges.
@@
1.1.3 Dynamic Packing and Covering.
Using the results thus far, the worst-case result for dynamic base packing, Theorem˜1.2, follows quite immediately: using Remark˜1.8, we know that the packing number can be expressed in terms of the ideal relative loads. We can approximate this by a greedy base collection, see Lemma˜1.6. And finally, we know how to maintain this under dynamic updates using Lemma˜1.7.
@@
The idea is to sample with $p=\frac{\log n}{\varepsilon^{2}\beta}$, such that the covering number in the sampled matroid becomes $\Theta(\log n/\varepsilon^{2})$ with high probability. The $\log n/\varepsilon^{2}$ stems from the use of Chernoff bounds. That means we are considering expressions of the form $\frac{|S|}{\operatorname{rk}(S)}$ under uniform sampling. Obviously, $|S|$ behaves as expected. The hard part is showing that $\operatorname{rk}(S)$ behaves well under sampling, i.e., does not change with the right choice of $p$. For graphic matroids, the expression simplifies by $\operatorname{rk}(S)=\#\text{vertices}$, which is clearly not affected by sampling. Hence these complications only occurs for general matroids.
@@
For the case of base packing a similar approach cannot be used. Again, the goal is to show how $\frac{|S|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{S})}$ behaves under uniform sampling, where $|S|$ behaves as expected. However, with the techniques known to us, it seems hard to show that $\operatorname{rk}(E)-\operatorname{rk}(\overline{S})$ behaves well with high probability. It remains an open question to obtain a dynamic base packing algorithm that updates independent of the packing number.

### v2

@@ 1 Introduction
One well-studied example of a matroid is the graphic matroid, where $E$ is some set of edges and $A\subseteq E$ is independent if and only if it is acyclic. Base packing corresponds to packing disjoint spanning trees and the base covering number corresponds to the arboricity of the graph. Throughout, we generalize results for the graphic matroid and we will state the algorithms for graphic matroids as a comparison.
@@
For dynamic matroids, we consider element insertions and deletions to the ground set. [^2] The decremental version (deletions only) of this has been studied in [BlikstadMNT23], and we argue below that this fully dynamic version is the natural and most general definition. When an element $e$ is deleted, the independent sets $\mathcal{I}$ simply restrict to the sets without $e$: $\{I\in\mathcal{I}:e\notin I\}$. In other words, $\mathcal{M}$ is restricted to $E\setminus\{e\}$. When an element $e$ is inserted, the adversary also decides on a collection $\mathcal{I}_{e}$ such that $(E\cup\{e\},\mathcal{I}\cup\mathcal{I}_{e})$ is a matroid and $e\in I$ for every $I\in\mathcal{I}_{e}$. This means that inserting and then deleting the same element results in the same matroid. The algorithm receives the element updates and can query the (new) independence sets.
@@
There exists a deterministic algorithm that, given a dynamic matroid $\mathcal{M}$ with weight function $w\colon E\to[1,\dots,W]$, maintains a minimum weight base, where each update uses $O(\log n)$ rank-queries.
Blikstad, Mukhopadhyay, Nanongkai, and Tu [BlikstadMNT23] provide an algorithm for maintaining a minimum weight basis under deletions. Each update requires $\tilde{O}(\sqrt{\operatorname{rk}(E)})$ worst-case rank-queries. However, using a ‘dynamic rank oracle’, they also ensure $\tilde{O}(\sqrt{\operatorname{rk}(E)})$ worst-case update time, as it can only answer queries, where the answer can be computed efficiently on a concrete matroid. The algorithm is based on the MST algorithm with $\tilde{O}(\sqrt{|V|})$ update time by Frederickson [Frederickson85], combined with the sparsification technique of Eppstein, Galil, Italiano, and Nissenzweig [EppsteinGIN97]. It seems likely that such a result can be extended to the fully dynamic setting. Blikstad et al. [BlikstadMNT23] use this as a subroutine and the authors seem to have optimized their result for their application in solving matroid union. The difference to our result, Proposition˜1.1, is that we have a much lower query time, but do not give guarantees on the update time, since the computation time for answering a query depends on the concrete matroid.
Both the algorithm [BlikstadMNT23] and our algorithm from LABEL:{lm:dyn_min_base} use a rank oracle. Sometimes, algorithms are developed using an independence oracle, which only provides whether $A\in\mathcal{I}$. Clearly, the rank oracle is stronger than the independence oracle. Statically, it is even known that it is strictly stronger: computing a minimum weight base can be done with $n$ simultaneous rank-queries, while this needs $\tilde{\Omega}(n^{1/3})$ rounds of simultaneous independence-queries [KarpUW88]. It would be interesting to see if there exists a decremental algorithm for maintaining a minimum weight base using a sublinear number of independence queries.
Minimum Weight Base in the Graphic Matroid.
@@
This is also known as matroid strength. The integral packing number is $\lfloor\Phi_{\mathcal{M}}\rfloor$. Edmonds [Edmonds65] showed that $\lfloor\Phi_{\mathcal{M}}\rfloor$ equals the number of disjoint bases that can be packed in $\mathcal{M}$.
@@
This is also known as matroid density. We define the integral covering number as $\lceil\beta_{\mathcal{M}}\rceil$. Edmonds [Edmonds65partition] showed that $\lceil\beta_{\mathcal{M}}\rceil$ equals the minimum number of bases necessary to cover $\mathcal{M}$. We omit the subscript if the matroid is clear from the context.
Computing exact packing and covering is a hard question. Even in certain specific matroids, like the graphic matroid (see the paragraph below) this is relatively slow. In this paper, we aim for efficient algorithms with $\operatorname{poly}\log n$ queries per update. In particular, we give the first dynamic algorithms for approximating the fractional packing number and the fractional covering number.
@@
Let $\varepsilon\in(0,1/2)$ be a parameter and let $\mathcal{M}$ be a dynamic matroid that at any moment contains at most $n$ elements. If $\Phi$ is upper bounded by $\Phi_{\max}$, we can deterministically maintain a $(1\pm\varepsilon)$-approximation of the fractional packing number with $O(\Phi_{\max}^{2}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ worst-case rank-queries per update or $O(\Phi_{\max}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ amortized rank-queries per update.
Since we give the first dynamic algorithm, our result can only be compared to using deterministic static algorithms to recompute the fractional packing number from scratch after every update. Using the state of the art by Chekuri and Quanrud [ChekuriQ17], we get $\tilde{O}(n\Phi_{\max}/\varepsilon^{2})$ queries per update. [^4] Chekuri and Quanrud [ChekuriQ17] only require the weaker independence-queries. However, even using a rank-oracle there is no known algorithm with $o(n)$ queries. The same holds for the results by [ChekuriQ17, Quanrud24] stated below. Hence, our algorithm provides an exponential improvement in terms of $n$.
@@
Let $\varepsilon\in(0,1/2)$ be a parameter and let $\mathcal{M}$ be a dynamic matroid that at any moment contains at most $n$ elements. If $\beta$ is upper bounded by $\beta_{\max}$, we can deterministically maintain a $(1\pm\varepsilon)$-approximation of the fractional covering number with $O(\beta_{\max}^{2}\cdot\varepsilon^{-4}\cdot\log^{3}n)$ worst-case rank-queries per update.
Again, there are no preexisting dynamic algorithms to compare our result to. Even static deterministic base covering is not as well studied. Quanrud [Quanrud24] conjectures that the techniques from Chekuri and Quanrud [ChekuriQ17] extend to deterministic approximate base covering in $\tilde{O}(n\beta_{\max}/\varepsilon^{2})$ queries, which would transfer to $\tilde{O}(n\beta_{\max}/\varepsilon^{2})$ queries per update.
@@
Once again, we compare to recomputing from scratch after every update using a static algorithm and obtain an exponential improvement. The state of the art for computing the static fractional covering number is by Quanrud [Quanrud24] and would result in $\tilde{O}(n+\operatorname{rk}(E)/\varepsilon^{3})$ queries per update. The same technique cannot be applied in the packing case to remove the dependence on $\Phi_{\text{max}}$, as we discuss in Section˜3.2.
Packing and Covering in the Graphic Matroid.
Dynamic tree-packing has been implicitly studied by Thorup [Thorup07]. This paper uses dynamic tree-packing due to its relation to min-cut. However, it implies a $(1\pm\varepsilon)$-approximation of the fractional tree packing number [^5] We realize that the naming conventions here overlap. Rather than disregarding the conventions completely, we write ‘tree-packing’ for the collection of trees and ‘tree packing’ when talking about the tree packing number. $\Phi\leq\Phi_{\max}$ in $\tilde{O}(\Phi_{\max}^{2}\log^{6}|E|/\varepsilon^{-4})$ amortized update time.
De Vos and Christiansen [deVosC24] give fully dynamic algorithms for $(1\pm\varepsilon)$-approximate arboricity with $O(\operatorname{poly}(\log|E|,\varepsilon^{-1}))$ update time against an adaptive adversary. Banerjee, Raman, and Saurabh [BanerjeeRS20] maintain the exact arboricity with $\tilde{O}(|E|)$ update time. As such, we believe that an algorithm with $\operatorname{poly}\log(|E|)$ queries per update for exact packing/covering for general matroids requires a breakthrough in techniques. We focus on obtaining $(1\pm\varepsilon)$-approximate results with $\operatorname{poly}\log(|E|)$ queries per update.
Other Applications.
The graphic matroid is one of the most studied matroids, where the above references show that dynamic base packing and covering has been studied directly. In other matroids, this is not the case. However, base packing and covering (and hence their dynamic versions) have other interesting applications. Here, we mention two: the linear matroid and the partition matroid, see e.g. [oxley92] for the definitions and connections.
In the linear matroid, base packing corresponds to the ability to decompose a matrix into many full-rank disjoint column sets. Base covering corresponds to finding multiple sets that form a basis for the spanned vector space, which has applications in (network) coding, see, e.g., [dougherty2007networks, casazza2013introduction]. In the partition matroid and generalizations thereof, base packing and covering correspond to multi-way assignment and various types of scheduling, see, e.g., [burkard1990constrained, kawase2021optimal].
Moreover, base packing and covering are highly connected to Shannon switching games [shannon1948mathematical, lehman1964solution].
@@ 1.1 Technical Overview
Our main technical tool are base collections, which we will introduce first. Then, we will sketch how we use them to obtain our packing and covering results, Theorems˜1.2, 1.3 and 1.4.
1.1.1 Base Collections
@@
Next, we want to define some ‘ideal’ relative load. Hereto, we first prove a corollary using Edmond’s theorem on base packing (Theorem˜4.1). We show that the lowest possible maximum load in base collections relate to the base packing number.
@@
We show that this is well-defined, meaning that the resulting ideal relative load values are independent of the concrete choices of the sets $A_{0}\subseteq E$ if there are multiple such sets that give $\frac{|A_{0}|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A_{0}})}=\Phi_{\mathcal{M}}$, in Section˜4.
@@
In the graphic matroid, it has been shown [Thorup07] that a greedy tree-packing $\mathcal{T}$ of size $\Theta(\Phi\log|V|/\varepsilon^{2})$ approximates the ideal packing well, resulting in every element having a relative load that differs from the ideal relative load by a small additive error as follows
@@
for all $e\in E$. We show a stronger statement for general matroids, for which we introduce the parameter $\gamma$ to provide a better trade-off between the error and the number of greedy bases. The parameter $\gamma$ satisfies $\Phi\leq\gamma\leq\beta$, which implies that $\Phi\leq\beta$ – which we have not shown yet. When considering the integer equivalents, the intuition is that $\lfloor\Phi\rfloor$ is the number of disjoint bases that fit in the matroid, and $\lceil\beta\rceil$ is the number of bases needed to cover the matroid, hence clearly $\lfloor\Phi\rfloor\leq\lceil\beta\rceil$. The fact that also $\phi\leq\beta$ is somewhat more technical, but easy to see using the structural results of Section˜1.1.2.
@@
The improvement over [Thorup07] in the special case of the graphic matroid can be seen as follows: Equation˜1 can be reformulated, by picking $\varepsilon$ accordingly, to obtain $|\ell^{\mathcal{B}}(e)-\ell^{*}(e)|\leq\varepsilon/\gamma$. However, it needs $|\mathcal{B}|=\Theta\left(\tfrac{\gamma^{2}}{\Phi}\log n/\varepsilon^{2}\right)$ bases. That is, it has a quadratic dependence on $\gamma$, where we obtain a linear dependence.
In other words, Thorup [Thorup07] showed that the values are an additive approximation, since the error is independent of $\ell^{*}(e)$. We show a multiplicative approximation, since the error depends linearly on $\ell^{*}(e)$. For our application, the latter leads to stronger results. To be precise, Lemma˜1.6 shows that a collection of at least $3\beta\log n/\varepsilon^{2}$ bases gives a $(1\pm\varepsilon)$-approximation of the covering number (using Theorem˜1.9). A generalization of Thorup’s version to matroids would give $|\mathcal{B}|=\Theta\left(\tfrac{\beta^{2}}{\Phi}\log n/\varepsilon^{2}\right)$.
@@
To the best of our knowledge, the community was not aware of this property, given in Lemma˜1.6, for the graphic matroid either. In particular, it means that we simplify the result of de Vos and Christiansen [deVosC24] for arboricity. They introduce an intricate procedure to artificially maintain $\Phi\approx\beta$ by adding virtual edges.
@@ 1.1.2 Structural Results
1.1.3 Dynamic Packing and Covering
Using the results thus far, the worst-case result for dynamic base packing, Theorem˜1.2, follows quite immediately: using Remark˜1.8, we know that the packing number can be expressed in terms of the ideal relative loads. We can approximate this by a greedy base collection, see Lemma˜1.6 and for more details refer to the proof of Theorem˜1.2 in Section˜3.1. And finally, we know how to maintain this under dynamic updates using Lemma˜1.7.
@@ 1.1.3 Dynamic Packing and Covering
Suppose we know (an approximation of) $\beta$. The idea is to sample with $p=\frac{\log n}{\varepsilon^{2}\beta}$, such that the covering number in the resulting sampled matroid will be $\Theta(\log n/\varepsilon^{2})$ with high probability. A constant approximation of $\beta$ suffices, which we can get from the approximation before the update. In total, we maintain $\log n$ copies of the algorithm: one for each possible estimate of $\beta$. The $\log n/\varepsilon^{2}$ term here stems from the use of Chernoff bounds.
Given this uniform sampling probability $p$, we consider the expressions $\frac{|S|}{\operatorname{rk}(S)}$ for each set $S$. Consider a set $S$ that maximizes this – the case of sets with smaller values is omitted in this overview for brevity, see Theorem˜1.4 for details. We show that their value will remain between $(1-\varepsilon)p\beta$ and $(1+\varepsilon)p\beta$ with high probability, where $\beta$ is the covering number in the original matroid. Obviously, the size of the set, $|S|$, behaves accordingly. The hard part is showing that the rank of the set, $\operatorname{rk}(S)$, behaves well under sampling, i.e., does not change by more than a factor $1\pm\varepsilon$ with our choice of $p$. For graphic matroids, the expression simplifies by $\operatorname{rk}(S)=\#\text{vertices in }S$, which is clearly not affected by sampling edges. Hence these complications only occur for general matroids.
@@
For the case of base packing a similar approach cannot be used. Again, the goal is to show how $\frac{|S|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{S})}$ behaves under uniform sampling. As before $|S|$ can be bounded by a Chernoff bound to be $(1\pm\varepsilon)p|S|$. However, with the techniques known to us, it seems hard to show that $\operatorname{rk}(E)-\operatorname{rk}(\overline{S})$ changes by at most a factor $(1\pm\varepsilon)$ with high probability. It remains an open question to obtain a dynamic base packing algorithm that updates independent of the packing number.
@@
Independent Work
The concurrent work of Arkhipov and Kolmogorov [arkhipov2025greedy] also obtains two of our results: Lemma˜1.6 for the special case that $\gamma=\beta$ and Theorem˜1.9. They use these results to show that they can maintain an approximation to the density of a graph by packing pseudoforests.